There is a population of 135 deer. Some of the deer in the population are albino which is a rare genetically inherited trait that is only expressed in the phenotype of the homozygous recessive individuals (aa). The phenotype of albinism is represented by a deficiency of skin and hair pogment melanin resulting in white hair and skin. In this population, 35 of the deer are albino and 7.5% of the 135 are heterozygous.

What are the expected genetyope frequencies if the population is not evolving? Show workd for full credit.

What is the chi squared statistic?
Answer #1

since 35 out of 135 deer are albino , therefore frequency of q^2 = 35/135 , i.e. 0.25

therefore q = 0.5 since p+q = 1, therefore p = 0.5

7.5% of 135 is 10 heterozygous individuals.

the expected genotype frequencies would be as follows:

p^2 = 0.25 , q^2 = 0.25 , 2pq = 0.5

the chi square statistic would be as follows:

Classes Observed (O) Expected (E) (O-E) (O-E)^2 (O-E)^2/E
p 0.5 0.5 0 0 0
q 0.5 0.5 0 0 0
Total 1 1 0 0 0

the total chi square observed is 0.

the degree of freedom is 2-1 = 1.

therefore, there is no significant difference between the two.